How do you calculate mass transferred by electroplating via electrolysis for:
Metal. Reduction half reactions
(Ag at 8.12A for 55 s). Ag+(aq) + e- —->Ag(s)
(Al at 20.5A for 105 s). Al3+(aq) +3e- ——>Al(s)
(Cu at 2.67 for 185 s). Cu2+(aq)+2e- —>Cu(s)
(Fe at 14.8 A for 75 s). Fe2+(aq)+2e- —->Fe(s)
( Ni at 25.0 A for 70 s). Ni2+(aq) +2e- —-> Ni(s)
1
Expert's answer
2020-05-10T09:21:47-0400

m=E x I x t/96500 (E - equivalent, I - current strength, t - time)

m= 108 x 8.12 x 55/96500=0.5 g of Ag.

m= 9 x 20.5 x 105/96500 = 0.2 g of Al.

m= 32 x 2.67 x 185 / 96500 = 0.164 g of Cu.

m= 28 x 14.8 x 75 /96500 = 0.322 g of Fe.

m= 29.5 x 25 x 70/96500 = 0.534 g of Ni.